# Foundations of Quantum Granular Computing with Effect-Based Granules, Algebraic Properties and Reference Architectures

### v1

@@ 4.2 Basic algebraic and probabilistic properties
The next result collects fundamental normalization and monotonicity properties of quantum granules.
@@
Let $\rho\in\mathcal{D}(\mathcal{H})$ and let $E,F$ be quantum granules on $\mathcal{H}$. Then:
@@
For (2), if $E\preceq F$ then $F-E\succeq 0$, and
@@ 4.3 Boolean islands of commuting granules
Let $\mathcal{E}=\{E_{1},\dots,E_{m}\}\subset\mathrm{Eff}(\mathcal{H})$ be a finite family of pairwise commuting quantum granules. Then there exists a finite probability space $(\Omega,\mathcal{F},\mu_{\rho})$, depending on $\rho$, such that:
@@
    each $E_{j}$ corresponds to a classical event $A_{j}\in\mathcal{F}$,
@@
    the algebra generated by $\{A_{j}\}_{j}$ is Boolean.
@@
Let $\mathcal{A}$ be the unital *-subalgebra of $\mathrm{L}(\mathcal{H})$ generated by $\mathcal{E}$. Since the effects $E_{j}$ commute pairwise, every element of $\mathcal{A}$ can be simultaneously diagonalized in some orthonormal basis of $\mathcal{H}$. In that basis, $\mathcal{A}$ is identified with the algebra of diagonal matrices, which is *-isomorphic to an algebra of functions on a finite set $\Omega$ [25, 10]. Each $E_{j}$ corresponds to a multiplication operator by a function $f_{j}:\Omega\to[0,1]$, and in the projective case, to a characteristic function $\mathbf{1}_{A_{j}}$ of some $A_{j}\subseteq\Omega$. The state $\rho$ induces a probability measure $\mu_{\rho}$ on $\Omega$, and $\mu_{\rho}(A_{j})=p_{\rho}(E_{j})$. The sets $\{A_{j}\}$ generate a Boolean algebra of events. ∎
In QGC terms, commuting granules form Boolean islands where classical granular reasoning is valid; incompatibility arises precisely when granules fail to commute. Theorem 2 shows that, within such an island, the granules $\{E_{j}\}$ can be identified with fuzzy-style sets $f_{j}:\Omega\to[0,1]$ under the classical probability measure $\mu_{\rho}$, such that all granular memberships and combinations reduce to standard probabilistic reasoning on a Boolean algebra.
@@ 4.4 Granular refinement under projective measurements
Then, for any quantum granule $E$,
$p_{\rho}(E)=\sum_{i}p_{i}\,p_{\rho_{i}}(E).$ (27)
@@
In QGC terms, Theorem 3 shows that the membership of a state $\rho$ in a coarse granule $E$ can be decomposed as a mixture of conditional memberships $p_{\rho_{i}}(E)$ given the outcomes of a PVM. This provides a quantum granular analogue of classical conditioning: before the measurement, $p_{\rho}(E)$ is a single global membership value, whereas after the measurement it is resolved into branches indexed by $i$, each weighted by $p_{i}$ and evaluated under the conditioned state $\rho_{i}$.
@@ 4.5 Evolution of quantum granules under channels
Let $\mathcal{E}:\mathrm{L}(\mathcal{H}_{\mathrm{in}})\to\mathrm{L}(\mathcal{H}_{\mathrm{out}})$ be a quantum channel, and let $E\in\mathrm{Eff}(\mathcal{H}_{\mathrm{out}})$ be a quantum granule on the output space. Then:
@@
    If $E\preceq F$, then $\mathcal{E}^{\dagger}(E)\preceq\mathcal{E}^{\dagger}(F)$,
@@
    For all $\rho\in\mathcal{D}(\mathcal{H}_{\mathrm{in}})$,
    $p_{\mathcal{E}(\rho)}(E)=p_{\rho}\bigl(\mathcal{E}^{\dagger}(E)\bigr).$ (28)
@@
In QGC terms, Theorem 4 shows that the effect of noise can be absorbed either into the state (Schrödinger picture) or into the granule (Heisenberg picture) without changing the resulting membership degrees. Operationally, measuring $E$ after a noisy evolution $\mathcal{E}$ is equivalent to measuring the dressed granule $\tilde{E}=\mathcal{E}^{\dagger}(E)$ on the ideal state $\rho$. This perspective will be important when considering NISQ-oriented implementations, where noise is unavoidable but can be modeled as a deformation of ideal granular structures.
@@ 4.6 Optimal decision granules
From a QGC perspective, Theorem 5 shows that $E^{\star}$ is an optimal quantum granule that partitions the space into “decide $\rho_{0}$” vs. “decide $\rho_{1}$” according to Bayes risk. Any decision rule of the form $\{E,I-E\}$ can be viewed as a two-outcome POVM, where $E$ represents an acceptance granule for hypothesis $\rho_{0}$ and IEI-E for $\rho_{1}$. The Helstrom granule $E^{\star}$ is optimal among all such effect-based rules, providing a canonical example of operator-valued granular decision-making and a template for fuzzy-like quantum decision schemes in QGC.
@@ 5 Quantum Granular Decision Systems (QGDS)
Quantum Granular Decision Systems (QGDS) provide a unified operational model for decision-making in Quantum Granular Computing. A QGDS integrates effect-based granules, measurement-induced partitions, and classical aggregation rules into a single inference pipeline. Conceptually, a QGDS generalizes classical granular decision systems—such as fuzzy or MFL-based decision architectures—by replacing set-theoretic granules with quantum effects and by allowing decisions to depend on non-commuting operators, contextual information, and quantum state representations.
A QGDS operates on quantum states encoded from data and processes them through a structured sequence of granular evaluations. Each evaluation is performed by an effect (quantum granule), and the resulting membership scores are combined into a final decision or inference. This mechanism extends classical soft-decision schemes to the operator domain and links them directly with optimality results from quantum detection theory (Sec. 4.6), where effect-based decision granules of Helstrom type provide canonical examples of Bayes-optimal quantum decisions.
@@ 5.1 Structure of a QGDS
A generic QGDS consists of four stages:
@@
    Classical or quantum input preprocessing. Data $x$ may be classical (numerical vectors, images, linguistic values) or already quantum (states prepared by upstream quantum processes). Classical inputs may be preprocessed via fuzzy sets, rough approximations, or MFL-based mediative granules to produce interpretable features.
@@
    Quantum encoding. The input is embedded into a state $\rho(x)$ on a suitable Hilbert space $\mathcal{H}$. The encoding may be amplitude-based, density-operator–based, or hybrid (e.g., encoding fuzzy memberships as populations in $\rho(x)$).
@@
    Granular evaluation via effects. A QGDS uses a family of effects $\{E_{j}\}_{j=1}^{m}$, typically arising from a PVM or, more generally, from a POVM. The output of this stage is the vector of granular probabilities
@@
    These values serve as soft quantum memberships, directly analogous to fuzzy degrees of membership but grounded in operator theory.
@@
    Common choices include margin-based rules, Bayesian optimal rules (e.g., Helstrom-type decision granules), or fuzzy-style aggregations.
Conceptually, the first two stages implement a data-to-state encoding, the third stage realizes effect-based granular evaluation as developed in Sec. 4, and the fourth stage applies a classical decision layer. This architecture matches the hybrid pipeline illustrated in Fig. 3, but now has a precise operational meaning grounded in quantum granulation and its connection to optimal quantum decision rules (Sec. 4.6).
@@ 5.2 QGDS and decision optimality
A central advantage of the QGDS framework is its compatibility with quantum-optimal decision granules. For binary decisions, Theorem 5 guarantees that the optimal granule is the projector onto the positive part of
@@
where $E^{\star}$ is the Helstrom granule. Thus, QGDS recovers fuzzy-like soft classification while retaining Bayes-optimality properties from quantum detection theory.
@@
In QGC terms, Theorem 5 provides a canonical example where the granular decision rule is not only soft and interpretable (via the memberships $\mu_{0},\mu_{1}$), but also provably optimal among all effect-based binary decisions. For multi-class problems, learned POVM-based QGDS architectures extend this principle by treating each effect $E_{j}$ as a class-specific decision granule, whose memberships are optimized directly in the operator domain.
@@ 5.3 QGDS pseudocode
The following pseudocode summarizes the hierarchical structure of a QGDS and is compatible with both simulated systems and noisy intermediate-scale quantum (NISQ) devices.
@@
Input: classical datum $x$ or prepared quantum state $\rho$
@@
1. Classical preprocessing (optional, if input is classical):
 If the input is a classical datum $x$, compute classical granules $G_{1},\dots,G_{k}$
and memberships $\mu_{i}(x)$ (e.g., fuzzy, rough, or MFL-based).
2. Quantum encoding:
 If the input is classical, map $x$ or $\{\mu_{i}(x)\}$ to a quantum state $\rho(x)\in\mathcal{D}(\mathcal{H})$.
If the input is already a prepared state, set $\rho(x):=\rho$.
@@ 4. Classical aggregation and decision:
 Compute $y=D(p_{1},\dots,p_{m})$ for a chosen rule $D$
@@
In practical terms, the QGDS algorithm can be read as follows. If the input is classical (e.g., a feature vector, an image, or a linguistic description), Step 1 optionally constructs classical information granules $G_{i}$ (for instance, fuzzy clusters or MFL-based mediative granules) and their memberships $\mu_{i}(x)$. These memberships provide an interpretable, granulation-based description of $x$ that can be fed into the quantum part of the pipeline. In Step 2, the system encodes either the raw datum $x$ or the granule-based features $\{\mu_{i}(x)\}$ into a quantum state $\rho(x)$ on a Hilbert space $\mathcal{H}$, for example, via amplitude encoding or density-operator encoding. If the input is already a prepared quantum state $\rho$ (as in a communication or sensing scenario), Step 1 is skipped, and Step 2 simply sets $\rho(x):=\rho$.
@@
Finally, Step 4 applies a classical aggregation rule $D$, which can range from a simple $\arg\max_{j}p_{j}(x)$ or a margin-based rule to a Bayes-optimal decision rule inspired by Theorem 5. The output $y$ is thus obtained by combining quantum granular memberships through a classical decision layer, completing the QGDS pipeline.
@@ 5.4 Interpretability and granule hierarchy
A notable feature of QGDS is the preservation of interpretability across quantum and classical stages. Classical granules (fuzzy sets, shadowed sets, MFL-based mediative granules) may feed into quantum granules through the encoding step, while quantum granular outputs $p_{j}$ may be passed to downstream fuzzy or decision-theoretic modules. This hierarchical granulation mirrors multi-level approximation structures in classical GrC, but is now enriched by non-commutative operations, contextual quantum information, and effect-based representations of uncertainty.
Through this synthesis, QGDS establishes a coherent computational model connecting operator-valued granules, quantum measurements, and classical inference. It forms the operational backbone upon which the case studies in Sec. 7 build, and provides a template for future architectures that integrate Quantum Granular Computing with broader quantum machine learning and decision-support systems.

### v2

@@ 4.2 Basic algebraic and probabilistic properties
We collect fundamental normalization and monotonicity properties of quantum granules, viewed through their membership functions.
@@
Let $\rho\in\mathcal{D}(\mathcal{H})$ and let $E,F\in\mathrm{Eff}(\mathcal{H})$ be effects (quantum granules) on $\mathcal{H}$. Then:
@@
which proves (1). For (2), if $E\preceq F$ then $F-E\succeq 0$, and hence
@@
Equivalently, for each granule $\mathcal{G}=(\mathcal{D}(\mathcal{H}),E,p_{E})$, Theorem 1 states that $p_{E}$ is a normalized, order-preserving membership map on the state space $\mathcal{D}(\mathcal{H})$.
@@
Classical coarse-graining of partition blocks is recovered by coarse-graining commuting projective granules.
Proposition 1 (Coarse-to-fine under commutation).
Let $\{P_{i}\}$ be pairwise orthogonal, commuting projectors with $\sum_{i}P_{i}=I$ (a PVM). For any index set $S$, define
$P_{S}:=\sum_{i\in S}P_{i}.$ (29)
Then $P_{S}$ is a projector that defines a coarser granule with membership
[…]
for all $\rho\in\mathcal{D}(\mathcal{H})$.
Proof.
Orthogonality and commutativity of the $P_{i}$ imply that $P_{S}$ is a projector. Linearity of the trace gives
[…]
Proposition 1 generalizes the coarse-to-fine behavior of classical partitions: a union of blocks corresponds to a coarser granule whose membership function is the sum of the memberships of the constituent blocks.
@@ 4.3 Boolean islands of commuting granules
Let $\mathcal{E}=\{E_{1},\dots,E_{m}\}\subset\mathrm{Eff}(\mathcal{H})$ be a finite family of pairwise commuting quantum granules. Then there exists a finite sample space $\Omega$, an algebra of subsets $\mathcal{F}$ of $\Omega$, and, for each state $\rho$, a probability measure $\mu_{\rho}$ on $(\Omega,\mathcal{F})$ such that:
@@
    each $E_{j}$ is represented by a classical function $f_{j}:\Omega\to[0,1]$,
@@
    the projection lattice generated by the spectral projectors of the $E_{j}$ is (via the above identification) isomorphic to a Boolean algebra of subsets of $\Omega$.
In the projective case, the $f_{j}$ reduce to characteristic functions $\mathbf{1}_{A_{j}}$ of events $A_{j}\in\mathcal{F}$, and $p_{\rho}(E_{j})=\mu_{\rho}(A_{j})$.
@@
Let $\mathcal{A}$ be the unital *-subalgebra of $\mathrm{L}(\mathcal{H})$ generated by $\mathcal{E}$. Since the effects $E_{j}$ commute pairwise, every element of $\mathcal{A}$ can be simultaneously diagonalized in some orthonormal basis of $\mathcal{H}$. In that basis, $\mathcal{A}$ is identified with the algebra of diagonal matrices, which is *-isomorphic to an algebra of functions on a finite set $\Omega$ [24, 11]. Each $E_{j}$ corresponds to a multiplication operator by a function $f_{j}:\Omega\to[0,1]$, and in the projective case, to a characteristic function $\mathbf{1}_{A_{j}}$ of some $A_{j}\subseteq\Omega$. The state $\rho$ induces a probability measure $\mu_{\rho}$ on $\Omega$, and $p_{\rho}(E_{j})=\mu_{\rho}(f_{j})$. The sets generated by the supports of the $f_{j}$ form a Boolean algebra of events. ∎
Viewed through QGC, commuting granules form Boolean islands where classical granular reasoning is valid; incompatibility arises precisely when granules fail to commute. Within such an island, the granules $\{E_{j}\}$ can be identified with fuzzy-style sets $f_{j}:\Omega\to[0,1]$ under the classical probability measure $\mu_{\rho}$ (for a fixed state $\rho$), so that granular memberships and their combinations reduce to standard probabilistic reasoning on a Boolean algebra.
@@ 4.4 Granular refinement under projective measurements
Let
$\rho^{\prime}\;:=\;\sum_{i}P_{i}\rho P_{i}\;=\;\sum_{i}p_{i}\,\rho_{i}$ (32)
denote the post-measurement state corresponding to a nonselective Lüders update. Then, for any effect $E\in\mathrm{Eff}(\mathcal{H})$ (quantum granule),
$p_{\rho^{\prime}}(E)\;=\;\sum_{i}p_{i}\,p_{\rho_{i}}(E).$ (33)
@@
Theorem 3 shows that the membership of the post-measurement state $\rho^{\prime}$ in a granule $E$ can be decomposed as a mixture of conditional memberships $p_{\rho_{i}}(E)$ given the outcomes of a PVM. This provides a quantum granular analogue of classical conditioning: after the measurement, $p_{\rho^{\prime}}(E)$ is resolved into branches indexed by $i$, each weighted by $p_{i}$ and evaluated under the conditioned state $\rho_{i}$.
@@ 4.5 Evolution of quantum granules under channels
Let $\mathcal{E}:\mathrm{L}(\mathcal{H}_{\mathrm{in}})\to\mathrm{L}(\mathcal{H}_{\mathrm{out}})$ be a quantum channel, and let $E\in\mathrm{Eff}(\mathcal{H}_{\mathrm{out}})$ be a quantum granule (effect) on the output space. Then:
@@
    If $E,F\in\mathrm{Eff}(\mathcal{H}_{\mathrm{out}})$ with $E\preceq F$, then $\mathcal{E}^{\dagger}(E)\preceq\mathcal{E}^{\dagger}(F)$,
@@
    3. For all $\rho\in\mathcal{D}(\mathcal{H}_{\mathrm{in}})$,
    $p_{\mathcal{E}(\rho)}(E)=p_{\rho}(\mathcal{E}^{\dagger}(E)).$ (34)
@@ 3. For all $\rho\in\mathcal{D}(\mathcal{H}_{\mathrm{in}})$,
Theorem 4 shows that the effect of noise can be absorbed either into the state (Schrödinger picture) or into the granule (Heisenberg picture) without changing the resulting membership degrees. Operationally, measuring $E$ after a noisy evolution $\mathcal{E}$ is equivalent to measuring the dressed granule $\tilde{E}=\mathcal{E}^{\dagger}(E)$ on the ideal state $\rho$. This perspective is important for NISQ-oriented implementations, where noise is unavoidable but can be modeled as a deformation of ideal granular structures.
@@ 4.6 Optimal decision granules
From the QGC perspective, Theorem 5 shows that $E^{\star}$ is an optimal quantum granule that partitions the space into “decide $\rho_{0}$” vs. “decide $\rho_{1}$” according to Bayes risk. Any decision rule of the form $\{E,I-E\}$ can be viewed as a two-outcome POVM, where $E$ represents an acceptance granule for hypothesis $\rho_{0}$ and IEI-E for $\rho_{1}$. The Helstrom granule $E^{\star}$ is optimal among all such effect-based rules, providing a canonical example of operator-valued granular decision-making and a template for fuzzy-like quantum decision schemes in QGC.
@@ 5 Quantum Granular Decision Systems (QGDS)
Quantum Granular Decision Systems (QGDS) provide an operational framework for decision-making in Quantum Granular Computing. A QGDS integrates effect-based granules, measurement-induced partitions, and classical aggregation rules into a single inference pipeline. It generalizes classical granular decision systems—such as fuzzy or MFL-based architectures—by replacing set-theoretic granules with quantum effects and by allowing decisions to depend on non-commuting operators, contextual information, and quantum state representations.
A QGDS operates on quantum states encoded from data and processes them through a prescribed sequence of granular evaluations. Each evaluation is specified by an effect (quantum granule), and the corresponding Born probabilities are interpreted as membership scores, which are subsequently aggregated into a final decision or inference. This construction extends classical soft-decision schemes to the operator domain and links them directly with optimality results from quantum detection theory (Sec. 4.6), where Helstrom-type decision granules provide canonical examples of Bayes-optimal quantum decisions.
@@ 5.1 Structure of a QGDS
A QGDS consists of four stages:
@@
    Classical or quantum input preprocessing. The input $x$ may be classical (numerical vectors, images, linguistic values) or already quantum (states prepared by upstream quantum processes). Classical data can be preprocessed via fuzzy sets, rough approximations, or MFL-based mediative granules to obtain structured, interpretable features.
@@
    Quantum encoding. The (preprocessed) input is embedded into a state $\rho(x)$ on a suitable Hilbert space $\mathcal{H}$. The encoding may be amplitude-based, density-operator–based, or hybrid (e.g., encoding fuzzy memberships as populations in $\rho(x)$).
@@
    Granular evaluation via effects. The system uses a family of effects $\{E_{j}\}_{j=1}^{m}$, typically arising from a PVM or, more generally, from a POVM. The output of this stage is the vector of granular probabilities
@@
    These values are interpreted as soft quantum memberships, analogous to fuzzy degrees of membership but grounded in operator theory.
@@
    Examples include margin-based rules, Bayesian optimal rules (e.g., Helstrom-type decision granules), or fuzzy-style aggregations.
The first two stages implement a data-to-state encoding, the third stage realizes effect-based granular evaluation as developed in Sec. 4, and the fourth stage applies a classical decision layer. This architecture matches the hybrid pipeline illustrated in Fig. 3, and provides a precise operational interpretation grounded in quantum granulation and its connection to optimal quantum decision rules (Sec. 4.6).
@@ 5.2 QGDS and decision optimality
A key advantage of the QGDS framework is its compatibility with quantum-optimal decision granules. For binary decisions, Theorem 5 guarantees that the optimal granule is the projector onto the positive part of
@@
where $E^{\star}$ is the Helstrom granule. In particular, QGDS recovers fuzzy-like soft classification while retaining Bayes-optimality properties from quantum detection theory.
@@
In the QGC framework, Theorem 5 provides a canonical example in which the granular decision rule is both soft and interpretable (via the memberships $\mu_{0},\mu_{1}$) and provably optimal among all effect-based binary decisions. For multi-class problems, learned POVM-based QGDS architectures extend this principle by treating each effect $E_{j}$ as a class-specific decision granule, whose memberships are optimized directly in the operator domain.
@@ 5.3 QGDS pseudocode
The following pseudocode summarizes the hierarchical structure of a QGDS and is compatible with both simulated systems and noisy intermediate-scale quantum (NISQ) devices. For clarity, the input may be either a classical datum $x$ or an already prepared quantum state $\rho_{\mathrm{in}}$; after Step 2, both cases are represented by a quantum state $\rho(x)$, and the subsequent stages are identical.
@@
Input: either a classical datum $x$ or a prepared quantum state $\rho_{\mathrm{in}}$
@@
1. Classical preprocessing (only if the input is classical):
if the input is a classical datum $x$ then
  compute classical granules $G_{1},\dots,G_{k}$ with memberships $\mu_{i}(x)$
(e.g., fuzzy, rough, or MFL-based).
2. State preparation / quantum encoding:
if the input is classical then
  map $x$ or $\{\mu_{i}(x)\}$ to a quantum state $\rho(x)\in D(H)$;
else (the input is already quantum) set $\rho(x):=\rho_{\mathrm{in}}$.
@@ 4. Classical aggregation and decision:
 Compute $y=D\bigl(p_{1}(x),\dots,p_{m}(x)\bigr)$ for a chosen rule $D$
@@
The QGDS algorithm can be interpreted as follows. When the input is classical (for example, a feature vector, an image, or a linguistic description), Step 1 optionally constructs classical information granules $G_{i}$ (such as fuzzy clusters or MFL-based mediative granules) with memberships $\mu_{i}(x)$. These memberships provide an interpretable, granulation-based description of $x$ that may be fed into the quantum part of the pipeline.
In Step 2, the system prepares the quantum state to be processed. If the input is classical, the raw datum $x$ or the granule-based features $\{\mu_{i}(x)\}$ are encoded into a quantum state $\rho(x)\in D(H)$ on a Hilbert space $H$, for instance via amplitude encoding or density-operator encoding. If the input is already a prepared quantum state $\rho_{\mathrm{in}}$ (as in a communication or sensing scenario), Step 1 is omitted and Step 2 simply sets $\rho(x):=\rho_{\mathrm{in}}$. Thus, from Step 3 onwards the classical-input and quantum-input modes share the same quantum granular evaluation and decision pipeline.
@@
Finally, Step 4 applies a classical aggregation rule $D$, ranging from a simple $\arg\max_{j}p_{j}(x)$ or a margin-based rule to a Bayes-optimal decision rule inspired by Theorem 5. The output $y$ is obtained by combining quantum granular memberships through this classical decision layer, completing the QGDS pipeline.
@@ 5.4 Interpretability and granule hierarchy
A notable feature of QGDS is the preservation of interpretability across quantum and classical stages. Classical granules (fuzzy sets, shadowed sets, MFL-based mediative granules) may feed into quantum granules through the encoding step, while quantum granular outputs $p_{j}(x)$ may be passed to downstream fuzzy or decision-theoretic modules. This hierarchical granulation mirrors multi-level approximation structures in classical GrC, but is now enriched by non-commutative operations, contextual quantum information, and effect-based representations of uncertainty.
Through this synthesis, QGDS provides a coherent computational model connecting operator-valued granules, quantum measurements, and classical inference. It also underpins the case studies in Sec. 7 and offers a template for architectures that integrate Quantum Granular Computing with broader quantum machine learning and decision-support systems.